# Statistical Inference for Pairwise Comparison Models

### v2

@@ 4 Examples
It can be verified that both conditions in (2) and (3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 2 (Thurstone–Mosteller model)
@@
where $\Phi(y)=\int_{-\infty}^{y}\varphi(z)\operatorname{d}\!{z}$ with $\varphi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. It can be verified that both conditions in (2) and (3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 3 (Rao–Kupper model)
@@
where $\theta>1$ is the threshold parameter which is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 4 (Davidson model)
@@
where $\theta>0$ is assumed to be prefixed. It can be verified that both conditions in (2) and (3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
Example 5 (Cumulative link model with four outcomes)
The cumulative link model is defined via the ordinalization of some distribution function $F$ (Agresti, 1992). The BT model, the Rao–Kupper model, and the Thurstone–Mosteller model can all be considered as specific cases of the cumulative link model with appropriate choices of $F$. In this example, we focus on the situation when $F$ is logistic with four outcomes, which will be used later in Section 5. In this case, $\mathbb{A}=\{-2,-1,1,2\}$ and the corresponding link function takes the form of
$f(1;y)=\frac{(\theta-1)e^{y}}{\left(\theta+e^{y}\right)\left(1+e^{y}\right)},\ \ \ f(2;y)=\frac{e^{y}}{\theta+e^{y}},$ (7)
where $\theta>1$ is the threshold parameter that is predetermined. It can be verified that both conditions in (2) and (3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$. The result above can be extended to the cumulative link model with the logistic link function and finite outcomes but involve more complicated formulas for asymptotic variances.
Example 6 (Paired cardinal model)
@@
where $\sigma>0$ is some fixed parameter. It can be verified that both conditions in (2) and (3) are satisfied if
@@
which is independent of ${\bm{u}}^{*}$. That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to N(0,1)$ as $n\to\infty$.
@@ 5.1 Synthetic data
We utilize the general random graph model introduced in Definition 2 to generate comparison graphs, where we set $p_{n}=n^{-1/2}$ and $q_{n}=p_{n}(\log n)$, with the individual edge probabilities chosen uniformly at random between them. The total number of subjects $n$ in all the tested models is chosen from the set $\{500,1000,2000\}$. To generate the utility vector ${\bm{u}}^{*}$, we uniformly select values from the range $[-M_{n},M_{n}]$, where the dynamic range $M_{n}$ is chosen from $\{1,\log(\log n)\}$. For global parameters in each model, we set $\theta=1$ in the Davidson model, $\theta=2$ in the Rao–Kupper model, and $\sigma=2$ in the paired cardinal model. Each simulation scenario is characterized by the pair $(n,M_{n})$ and repeated $300$ times. We report the average standard deviation of the estimated parameters and the coverage probability of the $95\%$ confidence interval, as shown in Table 2. We also examine the $z$-scores of the first coordinate of ${{\widehat{\bm{u}}}}$ over 300 experiments and plot the corresponding quantile-quantile plots against normal quantiles in Figure 1.
In Table 2, we observe a decreasing trend in the average standard deviation as $n$ increases, which is consistent with our theoretical findings on asymptotic variance in Equation (4). Additionally, we note variations in standard deviation relative to $M_{n}$ across different models. In particular, the standard deviations are increasing in $M_{n}$ in the Rao–Kupper and Davidson models, while remaining unchanged in the paired cardinal model. This observation aligns with their respective asymptotic variance formulas $\rho_{i}({\bm{u}}^{*})$ computed in Section 4; in particular, $\rho_{i}({\bm{u}}^{*})$ in the paired cardinal model is independent of $M_{n}$.
@@
5.2 Real data analysis
[…]
@@ 5.2 Real data analysis
To identify whether a significant distinction exists between the Big Three in tennis and Andy Murray, we employed a multiple testing procedure on the difference in their latent scores. Specifically, let $u_{1},u_{2},u_{3}$ and $u_{4}$ denote the latent score of Djokovic, Nadal, Federer, and Murray respectively. We consider the hypothesis testing with the null hypothesis $H_{0}:u_{i}=u_{4},i\in[3]$ and alternative hypothesis $H_{A}:$ there exists $i\in[3]$ such that $u_{i}\neq u_{4}$. The significance level $\alpha=0.05.$ We first obtained the maximum likelihood estimates of the latent scores $\widehat{{\bm{u}}}$, and the threshold parameter $\widehat{\theta}=2.32$. The estimator $\widehat{u}_{i}$ and estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$ of the 4 candidates are presented in Table 3. We construct the $z$-statistics $(\widehat{u}_{i}-\widehat{u}_{4})/{{(\widehat{\rho}_{i}+\widehat{\rho}_{4})^{1/2}}},\ i\in[3]$, which yields $p$-values $0.229$, $0.226$ and $0.334$ accordingly. For multiple testing, we adopt the Benjamini–Hochberg procedure (Benjamini and Hochberg, 1995). We order three $p$-values in ascending order and compare them with the corresponding critical values, that is $(\alpha)/3$, $(2\alpha)/3$, and $\alpha$. It turns out that all $p$-values exceed the corresponding critical value, implying that the difference in latent scores between Big Three and Murray is not statistically significant. Consequently, there is no sufficient statistical evidence under the selected threshold to reject the null hypothesis which supports the Big Four.
@@
Players Estimated scores $\widehat{{\bm{u}}}_{i}$ Estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$
@@ 6 Proof of Theorem 8
In this section, we provide a sketch of the proof for Theorem 8. Following a common route for establishing asymptotic normality, we apply Taylor’s expansion to the log-likelihood function and try to find the inverse of the corresponding Hessian matrix. However, the randomness of the Hessian matrix introduces additional challenges. The novelty of our approach lies in identifying the Hessian matrix as the weighted Laplacian matrix of appropriate random graphs, which can be carefully analyzed using truncated spectral methods.
@@
where the first inequality follows from the log-concavity of $f$ that is assumed under the validity assumption. Then the Hession of $l({\bm{u}})$, denoted by $H({\bm{u}})$ is given as
$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j.\end{array}\right.$
A crucial observation is that $H({\bm{u}})$ is the negation of an unnormalized graph Laplacian associated with a weighted graph on $[n]$ with edge weights $\partial_{ij}l({\bm{u}})$. Conditional on ${\mathcal{E}}$, $H({\bm{u}})$ is a random quantity due to the comparison outcomes $X_{ij}$. As such, we let $H^{*}({\bm{u}})={\mathbb{E}}[H({\bm{u}})\mid{\mathcal{E}}]$. It is easy to check that $-H^{*}({\bm{u}})$ remains an unnormalized weighted graph Laplacian that depends only on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and ${\bm{u}}$. In the following, we normalize $-H^{*}({\bm{u}}^{*})$ as
$-H^{*}({\bm{u}}^{*})={\mathcal{D}}^{1/2}{\mathcal{L}}_{sym}{\mathcal{D}}^{1/2}={\mathcal{D}}^{1/2}(I-{\mathcal{A}}){\mathcal{D}}^{1/2},$ (8)
where ${\mathcal{D}}=\text{diag}\{-H^{*}({\bm{u}}^{*})\}$ is the diagonal of $-H^{*}({\bm{u}}^{*})$ (degree matrix), ${\mathcal{L}}_{sym}$ is the normalized Laplacian of the weighted graph associated with $-H^{*}({\bm{u}}^{*})$, ${\mathcal{A}}=I-{\mathcal{L}}_{sym}$ is the normalized adjacency matrix, and $I$ is the identity matrix. By the mean value theorem,
@@
with ${\bm{W}}=({\bm{w}}_{1},\ldots,{\bm{w}}_{n}),$ and ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ (possibly distinct) points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$. It is worth noting that $\bar{H}({\bm{W}})$ is not a Hessian matrix (it is not even symmetric). Heuristically, thanks to the matrix concentration and uniform consistency of ${\widehat{\bm{u}}}$, $H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})=o(H({\bm{u}}^{*}))$ and $H({\bm{u}}^{*})-\bar{H}({\bm{W}})=o(H({\bm{u}}^{*}))$, where the $o(\cdot)$ is interpreted in an appropriate sense that will be made clear in Lemma 13.
With some rearrangement and substituting (8) into $H^{*}({\bm{u}}^{*})$, we have
${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (9)
@@
${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1},$ (10)
@@
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (11)
Based on the decomposition in (C.2), we will proceed to prove the desired asymptotic normality result as follows. We will show that the first term on the right-hand side of (C.2) appropriately normalizes $({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and gives the asymptotic normality. Meanwhile, through a meticulous analysis (Lemma 13), it can be established that the remaining terms are of order $o_{p}(1)$. Consequently, the desired follows by applying Slutsky’s theorem.
Lemma 13
@@
The proof of Lemma 13 is rather technical and deferred to the appendix. We now use the results in Lemma 13 to establish the asymptotic normality of ${\widehat{\bm{u}}}$. By statements (1) and (2) in Lemma 13 and Slutsky’s theorem, it remains to verify for every fixed $S\subset{\mathbb{N}}$ with $|S|=s$ ($S$ independent of $n$), $\{{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$. Since
${\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})={\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})+({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (12)
and $\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to N(0,I_{s})$ by the central limit theorem, the proof is finished if $\{({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}=o_{p}(1)$ componentwise. To verify this, we apply Chebyshev’s inequality. Note ${\mathbb{E}}[({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})]=0$ and the second moment calculation yields that for each $k\in[n]$ and the $k$th canonical basis vector $e_{k}\in{\mathbb{R}}^{n}$,

### v3

@@ 4 Examples
It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
Example 2 (Thurstone–Mosteller model).
@@
where $\Phi(y)=\int_{-\infty}^{y}\varphi(z)\operatorname{d}\!{z}$ with $\varphi(z)=(2\pi)^{-1/2}e^{-z^{2}/2}$. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
Remark 3.
Both the BT and Thurstone–Mosteller model models can be viewed as special cases of the Random Utility Model framework, which was formally developed for discrete choice analysis by mcfadden1973conditional. Specifically, the BT model arises from the assumption of i.i.d. Gumbel-distributed random utilities, while the Thurstone–Mosteller model assumes they are normally distributed. Our central findings can be generalized to the other types random utility models, which satisfies (A1)–(A5) in Definition 1.
Example 3 (Rao–Kupper model).
@@
where $\theta>1$ is the threshold parameter which is predetermined. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
Example 4 (Davidson model).
@@
where $\theta>0$ is assumed to be prefixed. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
Example 5 (Cumulative link model with four outcomes).
The cumulative link model is defined via the ordinalization of some distribution function $F$ (agresti1992analysis). The BT model, the Rao–Kupper model, and the Thurstone–Mosteller model can all be considered as specific cases of the cumulative link model with appropriate choices of $F$. In this example, we focus on the situation when $F$ is logistic with four outcomes, which will be used later in Section 5. In this case, $\mathbb{A}=\{-2,-1,1,2\}$ and the corresponding link function takes the form of
$f(1;y)=\frac{(\theta-1)e^{y}}{\left(\theta+e^{y}\right)\left(1+e^{y}\right)},\ \ \ f(2;y)=\frac{e^{y}}{\theta+e^{y}},$ (4.7)
where $\theta>1$ is the threshold parameter that is predetermined. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$. The result above can be extended to the cumulative link model with the logistic link function and finite outcomes but involve more complicated formulas for asymptotic variances.
Example 6 (Paired cardinal model).
@@
where $\sigma>0$ is assumed to be prefixed. It can be verified that both conditions in (3.2) and (3.3) are satisfied if
@@
which is independent of $\bm{u}$$^{*}$. That is, $\{{\rho_{i}({\bm{u}}^{*})}\}^{-1/2}(\widehat{u}_{i}-u^{*}_{i})\to\mathcal{N}(0,1)$ as $n\to\infty$.
@@ 5.1 Synthetic data
We utilize the general random graph model introduced in Definition 2 to generate comparison graphs, where we set $p_{n}=n^{-1/2}$ and $q_{n}=p_{n}(\log n)$, with the individual edge probabilities chosen uniformly at random between them. The total number of subjects $n$ in all the tested models is chosen from the set $\{500,1000,2000\}$. To generate the utility vector $\bm{u}$$^{*}$, we uniformly select values from the range $M_{n},M_{n}]$, where the dynamic range $M_{n}$ is chosen from $\{1,\log(\log n)\}$. For global parameters in each model, we set $\theta=1$ in the Davidson model, $\theta=2$ in the Rao–Kupper model, and $\sigma=2$ in the paired cardinal model. Each simulation scenario is characterized by the pair $(n,M_{n})$ and repeated $300$ times. We report the average standard deviation of the estimated parameters and the coverage probability of the $95\%$ confidence interval, as shown in Table 2. We also examine the $z$-scores of the first coordinate of $^{{\widehat{\bm{u}}}}$ over 300 experiments and plot the corresponding quantile-quantile plots against normal quantiles in Figure 1.
In Table 2, we observe a decreasing trend in the average standard deviation as $n$ increases, which is consistent with our theoretical findings on asymptotic variance in (3.4). Additionally, we note variations in standard deviation relative to $M_{n}$ across different models. In particular, the standard deviations are increasing in $M_{n}$ in the Rao–Kupper and Davidson models, while remaining unchanged in the paired cardinal model. This observation aligns with their respective asymptotic variance formulas $\rho_{i}({\bm{u}}^{*})$ computed in Section 4; in particular, $\rho_{i}({\bm{u}}^{*})$ in the paired cardinal model is independent of $M_{n}$.
@@
5.2 Real data example
In the realm of men’s tennis, the dominance of Novak Djokovic, Roger Federer, Rafael Nadal, and Andy Murray, collectively known as the Big Four, has undeniably dominated the first quarter of the 21st century. However, recent discussions have emerged, advocating for the exclusion of Andy Murray in what could potentially become the Big Three. In this section, we consider a statistical approach to model the ATP data and uncover the latent scores that provide evidence in favour of the Big Four. We consider modeling the comparison data from ATP111www.tennis-data.co.uk using the cumulative link model with four outcomes from Example 5. The ATP dataset collects the comparison results of each game from 1968 to 2016 by 2742 players. In this study, we concentrate on the best-of-3 matches, which result in four possible outcomes $\{\text{0:2, 1:2, 2:1, 2:0}\}$, represented as ${\mathbb{A}}=\{-2,-1,1,2\}.$ In particular, transferring the ordinal outcomes (4.7) into the current setting reads
@@ 5.2 Real data example
The estimated latent scores $\widehat{u}_{i}$ and corresponding standard deviation $(\widehat{\rho}_{i})^{1/2}$ of the 4 candidates are presented in Table 3, and it shows that Djokovic, Nadal, and Federer exhibit comparable performance levels, while Murray’s score is slightly lower. Moreover, the overlap in standard deviations implies potential ranking uncertainty.
To quantify the ranking uncertainty, we use parametric bootstrap to simulate 10,000 samples by the asymptotic normal distributions $\mathcal{N}(\widehat{u}_{i},\widehat{\rho}_{i})$. Then we calculate empirical ranks across Monte Carlo samples and construct histogram for each rank position, shown in Figure 2(a), where each bar indicates the probability of each player ranking from 1st through 4th. Figure 2(a) reveals that Djokovic has the highest probability of achieving the top rank, closely followed by Nadal and Federer, while Murray consistently ranks last, indicating the gap between him and the Big Three.
To further investigate pairwise competitiveness among players, we construct a posterior probability matrix. It summarizes the likelihood that player $i$ outperforms player $j$ based on their estimates. Specifically, we compute:
${\mathbb{P}}\left(u_{i}>u_{j}\right)=\mathbb{P}\left(Z_{i}-Z_{j}>0\right),$
where $Z_{i}\sim\mathcal{N}\left(\widehat{u}_{i},\widehat{\rho}_{i}\right)$ and $Z_{j}\sim\mathcal{N}\left(\widehat{u}_{j},\widehat{\rho}_{j}\right)$ are independently drawn from the posterior distributions of players $i$ and $j$, respectively. The asymptotic distribution guarantees that
$\mathbb{P}\left(u_{i}>u_{j}\right)=\Phi\left(\frac{\widehat{u}_{i}-\widehat{u}_{j}}{\sqrt{\widehat{\rho}_{i}+\widehat{\rho}_{j}}}\right),$
where $\Phi(\cdot)$ denotes the standard normal cumulative distribution function. The heatmap in Figure 2(b) displays the resulting posterior win probability matrix , where each entry $(i,j)$ shows the empirical probability that player $i$ wins player $j$ in terms of latent score under 10,000 repetitions. We can observe that Djokovic and Nadal have approximately an equal chance of outperforming each other, and both have $60$-$65\%$ probability of outperforming Federer. Murray has the lowest win probabilities against all other players, consistent with his lower latent score. In general, Djokovic and Nadal are the top players with extremely close latent strength, Federer is a strong challenger but slightly behind and Murray most frequently ranked the lowest.
To further assess the uncertainty in individual scores, we construct confidence intervals for latent scores. Figure 3(a) presents the $95\%$ confidence intervals for $\widehat{u}_{i}$ of each player under the CLM4 model. We use standard normal approximation to the asymptotic distribution to construct the confidence intervals, i.e., $\widehat{u}_{i}\pm z_{0.95}\sqrt{\widehat{\rho}_{i}}$. Figure 3(a) confirms the dominance of Djokovic and Nadal, the competitive strength of Federer, and a notable performance gap for Murray. However, the overlapping of all confidence intervals shows that these 4 players are not significantly different. To further validate the distinction with statistically significance, we construct the $95\%$ confidence interval for latent score difference. To construct the intervals, we apply the Bonferroni-adjusted critical value to the pairwise normal differences
$\widehat{u}_{i}-\widehat{u}_{j}\pm z_{1-\alpha/(2K)}\cdot\sqrt{\widehat{\rho}_{i}+\widehat{\rho}_{j}}$
where $K$ is the number of pairwise comparisons and $z_{1-\alpha/(2K)}$ is the Bonferroni-corrected quantile. Figure 3(b) shows that none of the intervals exclude point zero, indicating that although Murray seems rank the least from previous experiments, all four players are not significantly different from others at the $95\%$ confidence level.
In summary, both point estimates and posterior analyses indicate a discernible ranking between Murray and the Big Three. However, the simultaneous confidence intervals reveal that none of the pairwise differences are statistically significant. This highlights the critical role of formal statistical inference in practical applications.
@@
Players Estimated scores $\widehat{u}_{i}$ Estimated standard deviation $(\widehat{\rho}_{i})^{1/2}$
@@
(a) Posterior ranking probability histogram.
(b) Win probability heatmap.
Figure 2: The posterior ranking and pairwise win probability for each player.
(a) The $95\%$ confidence interval of latent score.
(b) The $95\%$ confidence interval of difference.
Figure 3: The $95\%$ confidence interval of the latent score of individual players and the pairwise difference between players’ scores.
@@ 6 Proof of Theorem 1
In this section, we provide a sketch of the proof for Theorem 1. Following a common route for establishing asymptotic normality, we apply Taylor’s expansion to the log-likelihood function and try to find the inverse of the corresponding Hessian matrix. However, the randomness of the Hessian matrix introduces additional challenges. The crux of our approach lies in identifying the corresponding Fisher information matrix as the weighted Laplacian matrix of appropriate random graphs, which can be carefully analyzed using truncated spectral methods. Although leveraging the Fisher information matrix to derive asymptotic normality in network models is well established (MR1724040; yan2013central; han2023unified), existing results require specific model structures, particularly those with logistic-type link functions. To the best of our knowledge, our work provides the first general framework that maintains optimality conditions while accommodating broader model specifications.
@@
where the first inequality follows from the log-concavity of $f$ that is assumed under the validity assumption. Then the Hessian of $l({\bm{u}})$, denoted by $H({\bm{u}})$ is given as
$\{H({\bm{u}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j.\end{array}\right$.
A crucial observation is that $H({\bm{u}})$ is the negation of an unnormalized graph Laplacian associated with a weighted graph on $[n]$ with edge weights $\partial_{ij}l({\bm{u}})$. Conditional on $\mathcal{E}$}, $H({\bm{u}})$ is a random quantity due to the comparison outcomes $X_{ij}$. As such, we let $H^{*}({\bm{u}})={\mathbb{E}}[H({\bm{u}})\mid{\mathcal{E}}]$. It is easy to check that the Fisher information matrix, $H^{*}({\bm{u}})$ remains an unnormalized weighted graph Laplacian that depends only on ${\mathcal{G}}({\mathcal{V}},{\mathcal{E}})$ and $\bm{u}$}. In the following, we normalize $H^{*}({\bm{u}}^{*})$ as
$-H^{*}({\bm{u}}^{*})={\mathcal{D}}^{1/2}{\mathcal{L}}_{sym}{\mathcal{D}}^{1/2}={\mathcal{D}}^{1/2}(I-{\mathcal{A}}){\mathcal{D}}^{1/2},$ (6.8)
where ${\mathcal{D}}=\text{diag}\{-H^{*}({\bm{u}}^{*})\}$ is the diagonal of $H^{*}({\bm{u}}^{*})$ (degree matrix), ${\mathcal{L}}_{sym}$ is the normalized Laplacian of the weighted graph associated with $H^{*}({\bm{u}}^{*})$, ${\mathcal{A}}=I-{\mathcal{L}}_{sym}$ is the normalized adjacency matrix, and $I$ is the identity matrix. By the mean value theorem,
@@
with ${\bm{W}}=({\bm{w}}_{1},\ldots,{\bm{w}}_{n}),$ and ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ (possibly distinct) points on the line segment between $^{\widehat{\bm{u}}}$ and $\bm{u}$$^{*}$. It is worth noting that $\bar{H}({\bm{W}})$ is not a Hessian matrix (it is not even symmetric). Heuristically, thanks to the matrix concentration and uniform consistency of $^{\widehat{\bm{u}}}$, $H^{*}({\bm{u}}^{*})-H({\bm{u}}^{*})=o(H({\bm{u}}^{*}))$ and $H({\bm{u}}^{*})-\bar{H}({\bm{W}})=o(H({\bm{u}}^{*}))$, where the $o(\cdot)$ is interpreted in an appropriate sense that will be made clear in Lemma 1.
With some rearrangement and substituting (6.8) into $H^{*}({\bm{u}}^{*})$, we have
${\mathcal{D}}^{1/2}({\widehat{\bm{u}}}-{\bm{u}}^{*})$ $=-{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (6.9)
@@
${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1},$ (6.10)
@@
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (6.11)
Based on the decomposition in (6.9), we will proceed to prove the desired asymptotic normality result as follows. We will show that the first term on the right-hand side of (6.9) appropriately normalizes $({\widehat{\bm{u}}}-{\bm{u}}^{*})$ and gives the asymptotic normality. Meanwhile, through a meticulous analysis (Lemma 1), it can be established that the remaining terms are of order $o_{p}(1)$. Consequently, the desired follows by applying Slutsky’s theorem.
Lemma 1.
@@
The proof of Lemma 1 is rather technical and deferred to the appendix. We now use the results in Lemma 1 to establish the asymptotic normality of $^{\widehat{\bm{u}}}$. By statements (1) and (2) in Lemma 1 and Slutsky’s theorem, it remains to verify for every fixed $S\subset{\mathbb{N}}$ with $|S|=s$ ($S$ independent of $n$), $\{{\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to\mathcal{N}(0,I_{s})$. Since
${\mathcal{L}}^{\dagger}_{sym}{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})={\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})+({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})$ (6.12)
and $\{{\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}\to\mathcal{N}(0,I_{s})$ by the central limit theorem, the proof is finished if $\{({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})\}_{S}=o_{p}(1)$ componentwise. To verify this, we apply Chebyshev’s inequality. Note ${\mathbb{E}}[({\mathcal{L}}^{\dagger}_{sym}-I){\mathcal{D}}^{-1/2}\nabla l({\bm{u}}^{*})]=0$ and the second moment calculation yields that for each $k\in[n]$ and the $k$th canonical basis vector $e_{k}\in{\mathbb{R}}^{n}$,